# Off the Planckian Locus: Using 2D Chromaticity to Improve In-Camera Color

### v2

@@ 4.2 Lightbox Dataset Results
Oracle 1D 0.00 0.01 5.37 2.91 4.70 7.27 9.49 5.83 3.22 5.20 7.85 10.17 7.07 4.55 6.41 8.62 12.14 6.51 4.09 5.94 8.20 11.06
2-CST 1D 0.07 0.01 7.21 3.78 6.06 9.02 13.18 8.15 4.41 6.63 10.23 15.08 9.18 5.68 7.91 11.29 15.74 8.61 5.28 7.64 10.66 14.76
3-CST 1D 0.11 0.01 7.20 3.77 6.02 9.02 13.19 8.13 4.40 6.65 10.17 15.10 9.16 5.68 7.93 11.28 15.65 8.62 5.28 7.65 10.67 14.76
NN 1D 15.43 0.01 8.04 4.05 6.64 10.22 15.00 9.21 4.57 7.20 11.62 17.90 9.62 5.60 8.21 11.97 17.20 9.48 5.34 8.12 11.54 16.90
CCCNN 1D 68.51 13.39 7.03 3.54 6.12 9.81 12.55 7.72 4.11 6.74 10.49 13.85 8.30 4.92 7.27 10.65 14.15 8.27 4.53 7.74 10.63 14.05
EXPINV 1D 68.53 13.39 6.94 3.53 5.97 9.57 12.29 7.64 3.76 6.67 10.38 14.01 8.18 4.87 7.26 10.39 13.96 8.22 4.83 7.42 10.24 14.02
CST-MLP 1D 1.41 0.01 6.93 3.55 5.99 8.97 12.60 7.70 3.95 6.53 9.71 14.55 8.35 4.90 7.51 10.81 14.66 8.06 4.46 7.39 10.22 14.10
NN 2D 16.97 0.01 7.17 3.60 5.96 9.03 12.92 7.79 3.94 6.37 9.81 14.30 8.75 5.27 7.42 10.69 15.21 8.37 4.80 7.16 10.14 14.71
CCCNN 2D 69.01 13.48 6.63 3.69 5.62 9.17 11.40 6.78 3.58 6.07 9.27 11.61 8.07 4.88 7.42 10.33 12.91 7.76 4.86 7.66 9.47 11.96
EXPINV 2D 69.03 13.48 6.48 3.45 5.84 8.98 11.03 7.07 3.98 6.52 9.39 11.93 7.77 4.74 7.10 10.02 12.63 7.68 4.37 7.51 9.72 12.20
CST-MLP 2D 1.54 0.01 6.60 3.50 5.95 8.60 11.49 7.15 3.80 6.31 9.17 12.85 7.91 4.66 7.25 10.28 13.27 7.70 4.26 7.28 9.88 13.23
Table 1: Quantitative evaluation on lightbox dataset. We compare angular error and $\Delta E_{2000}$ of traditional 2-CST/3-CST interpolation [AdobeDNG, Hakki_CVPR18], NN [nearestneighbors], CCCNN [cccnn], EXPINV [expinv], and our CST-MLP. For NN, CCCNN, EXPINV, and CST-MLP, we evaluate both 1D CCT-based implementations and our proposed 2D chromaticity-based adaptations (highlighted in blue).
Comparing CST-MLP with per-pixel MLPs.
We find our 2D CST-MLP outperforms 1D MLP-based direct prediction baselines, CCCNN [cccnn] and EXPINV [expinv]. Additionally, highlighting the strength of utilizing 2D chromaticity input, our 2D chromaticity-based variants of CCCNN [cccnn] and EXPINV [expinv] achieve lower $\Delta E_{2000}$ in many cases and surpass CST-MLP under challenging illumination conditions (higher percentiles). This performance gain arises because per-pixel MLPs can model more complex mappings, whereas our method is limited to predicting a linear transformation. However, per-pixel methods remain unsuitable for two reasons. First, CST-MLP has a footprint of 1.5 KB and 0.01 MMACs, compared to 69KB and 13 MMACs for per-pixel methods. CST-MLP is efficient because it predicts a single global transform. In Figure 6, we show that the computational cost of a per-pixel MLP far exceeds our CST-MLP at various input resolutions. Second, as shown in Figure 7, applying per-pixel MLP methods to full-size images can produce color casts, likely because each pixel is transformed independently, leading to spatial inconsistencies. In contrast, CST-MLP applies a single transformation across the image, ensuring spatial consistency.
In-the-wild results.
@@
After calibrating the CST prediction methods with lightbox images, we evaluate on a smaller set of 64 in-the-wild photos captured with both smartphones (Pixel and Samsung). As shown in Figure 7, methods that rely on 2D chromaticity deliver qualitatively superior results, enabling more accurate color correction under challenging LED illumination across diverse regions containing skin tones, posters, and craft materials. We also observe that EXPINV [expinv], a per-pixel prediction method, exhibits color casts when applied to full-size images. Additional comparisons are provided in Supplementary Section S2.
Ablation.
In Table 6, we ablate our choice of using a 2D xy chromaticity coordinate against modeling the illuminant in 1D CCT space or in 2D raw space. Additionally, we find that our noise augmentation strategy is crucial for our method to generalize to out-of-domain illuminant chromaticities. We further ablate CST-MLP with varying layer and node counts in Supplementary Section S1.4.
@@
Figure 5: Computational cost comparison. We plot the MACs (log scale) at different resolutions for 2-CST [AdobeDNG] interpolation, EXPINV [expinv], and 2D CST-MLP [expinv]. Our CST-MLP has nearly the same computational cost as traditional methods, since the CST is computed once and applied to the entire image. In contrast, approaches such as EXPINV must evaluate the MLP at every pixel.
Input [-1pt]coord. Noise Aug. Lightbox In-the-wild
@@
1D (CCT) 4.03 8.35 4.86 8.48
2D (raw) 3.58 7.99 5.39 9.64
2D (raw) 3.70 7.91 4.92 8.79
2D (xy) 3.53 7.92 5.04 9.12
2D (xy) 3.57 7.91 4.47 8.29
Figure 6: Ablation of CST-MLP. We evaluate CST-MLP on Pixel camera data from both lightbox and in-the-wild scenes, testing 1D and 2D input coordinates with and without noise augmentation.
@@
Figure 8: Colorimetric mapping results on spatially-varying multi-illuminant data from LSMI [lsmi]. We visualize two multi-illuminant scenes from the Galaxy camera. Row structure for each half: Top row displays raw-RGB image, white-balanced raw-RGB image, and display-referred renderings; middle row shows contrast-stretched illumination map, first raw-RGB color chart, and display-referred renderings; bottom row presents blending map, second raw-RGB color chart, and display-referred renderings. Methods from left to right: 2-CST [AdobeDNG], 1D NN [nearestneighbors], 1D CST-MLP, 2D NN [nearestneighbors], 2D CST-MLP, and Oracle. Angular errors are displayed at the bottom left of each color chart image. Key findings: (1) Our 2D chromaticity-based methods can be extended to multi-illuminant scenes with knowledge of the illumination map and per-pixel blending maps. (2) Our 2D methods continue to exceed the performance of 1D counterparts.
@@ 4.3 LSMI Dataset Results
Finally, we evaluate CST-MLP on single-illuminant scenes and further extend it to mixed-illuminant scenes, both sourced from the LSMI [lsmi] dataset.
Single-illuminant results.
We train our CST-MLP and all baseline models on single-illuminant images from the LSMI dataset, since we do not have access to lightbox captures. We filter the LSMI dataset to include scenes where all patches on all color charts are not clipped, resulting in 1504 images for Galaxy and 1414 for Sony. We then split the data into train, validation, and test sets using 50%, 20%, and 30% of the images, respectively. For this method, we are unable to compare with CCCNN [cccnn] and EXPINV [expinv] because these methods require access to the cameras’ spectral sensitivities, which are not provided with the LSMI dataset. Summarized results comparing our 2D CST-MLP with 2-CST interpolation (using DNG CST matrices), nearest-neighbor (NN), and 1D CST-MLP baselines are presented in Table 2 (top). We find that our 2D CST-MLP achieves the highest performance on this dataset. Additional details and results are provided in Supplementary Section S3.
Single-illuminant
Method Dim Galaxy Sony
[…]
Oracle 1D 2.38 4.90 1.59 3.52
2-CST [AdobeDNG] 1D 5.00 8.43 8.40 9.80
NN [nearestneighbors] 1D 4.46 7.40 3.07 5.19
CST-MLP 1D 3.77 6.30 2.70 4.55
NN [nearestneighbors] 2D 3.98 6.56 2.48 4.39
CST-MLP 2D 3.49 5.82 2.49 4.22
Multi-illuminant
Oracle 1D 2.06 3.92 1.44 3.01
2-CST [AdobeDNG] 1D 3.88 6.87 8.14 9.08
NN [nearestneighbors] 1D 3.23 5.42 2.43 4.26
CST-MLP 1D 2.52 4.28 1.88 3.40
NN [nearestneighbors] 2D 2.83 4.76 1.93 3.61
CST-MLP 2D 2.26 3.88 1.82 3.28
Table 2: LSMI results under single- and mixed-illuminant settings. We compare our 2D chromaticity-based CST-MLP and NN [nearestneighbors] (highlighted blue) with baselines on Galaxy and Sony data from LSMI [lsmi], reporting both angular error and $\Delta E_{2000}$.
Multi-illuminant results.
Assuming access to per-pixel illuminant information and blending maps that quantify each illuminant’s contribution, we can apply our MLP trained on single-illuminant images to multi-illuminant scenes. Specifically, for each illuminant, we predict and apply the CST to the white-balanced image to generate multiple corrected versions. We then merge these corrected images using the blending map as weights. A detailed description of this procedure is provided in Supplementary Section S3. In Table 2 (bottom), we present quantitative comparisons of our method. Additionally, in Figure 8, we present qualitative results on 2- and 3-illuminant scenes from the LSMI dataset. We observe that our 2D CST-MLP outperforms other methods quantitatively and qualitatively, demonstrating the benefits of using a 2D chromaticity representation in multi-illuminant scenarios. Additional details and results are provided in Supplementary Section S3.

### v3

@@ 4.1 Evaluation Details
Oracle 1D 0.00 0.01 5.37 2.91 4.70 9.49 5.83 3.22 5.20 10.17 7.07 4.55 6.41 12.14 6.51 4.09 5.94 11.06
2-CST [AdobeDNG] 1D 0.07 0.01 7.21 3.78 6.06 13.18 8.15 4.41 6.63 15.08 9.18 5.68 7.91 15.74 8.61 5.28 7.64 14.76
3-CST [Hakki_CVPR18] 1D 0.11 0.01 7.20 3.77 6.02 13.19 8.13 4.40 6.65 15.10 9.16 5.68 7.93 15.65 8.62 5.28 7.65 14.76
NN [nearestneighbors] 1D 15.43 0.01 8.02 4.00 6.63 15.03 9.21 4.57 7.20 17.91 9.53 5.57 8.07 17.03 9.49 5.36 8.18 16.81
CCCNN [cccnn] 1D 68.51 13.39 7.00 3.53 5.90 12.55 7.93 4.16 6.53 14.42 8.36 4.91 7.40 14.39 8.45 5.05 7.96 13.93
EXPINV [expinv] 1D 68.53 13.39 7.03 3.55 5.98 12.42 7.76 3.99 6.75 14.25 8.09 4.83 7.14 13.97 8.17 5.07 7.29 13.98
CST-MLP 1D 1.41 0.01 7.09 3.86 6.20 12.56 7.81 4.25 6.87 14.35 8.48 4.99 7.45 14.58 8.14 4.65 7.51 14.10
CST-MLP(RP) 1D 2.57 0.01 7.60 4.97 6.83 12.60 8.28 5.36 7.25 14.01 8.19 4.60 6.89 14.82 8.39 4.84 7.09 14.65
NN [nearestneighbors] 2D 16.97 0.01 7.29 3.60 5.99 13.17 7.79 3.92 6.37 14.21 8.74 5.26 7.35 15.30 8.37 4.80 7.16 14.71
CCCNN [cccnn] 2D 69.01 13.48 6.63 3.69 5.62 11.40 6.78 3.58 6.07 11.61 8.07 4.88 7.42 12.91 7.76 4.86 7.66 11.96
EXPINV [expinv] 2D 69.03 13.48 6.48 3.45 5.84 11.03 7.07 3.98 6.52 11.93 7.77 4.74 7.10 12.63 7.68 4.37 7.51 12.20
CST-MLP 2D 1.54 0.01 6.73 3.66 6.07 11.58 7.17 3.96 6.41 12.55 7.95 4.78 7.21 13.26 7.74 4.38 7.29 13.11
CST-MLP(RP) 2D 2.70 0.01 7.16 4.57 6.48 11.56 7.47 4.28 6.83 12.68 7.74 4.50 6.64 13.84 8.08 4.56 7.10 13.97
Noise augmentation. When queried with out-of-domain chromaticities not seen in our training set, our 2D CST-MLP can sometimes fail to generalize. As illustrated in  Fig.˜4, the training distribution doesn’t contain all in-the-wild illuminants. We address this issue by adding Gaussian noise to the input chromaticity coordinates during training [ling2025stochastic], enabling models to observe chromaticities that may be out of the training distribution. For CST-MLP, we add noise with standard deviation $\sigma_{0}=0.05$ to the normalized input coordinates for all training iterations. We show the positive effects of this augmentation in our ablation study (Tab.˜2). Finally, we exclude this augmentation for direct-prediction MLP baselines (EXPINV [expinv], CCCNN [cccnn]) as it can degrade their performance (see Supplementary Sec.˜S1.3).
Metrics. For evaluation, we utilize angular error following previous work [Hakki_CVPR18]. We also compare $\Delta E_{2000}$ after applying a uniform scaling to all colors to match the Y channel of the third gray patch (bottom row of color chart). Finally, we report model parameter sizes (in KB) and inference costs (in MACs) when applied to a $1080\times 720$ image.
4.2 Lightbox Dataset Results
We compare all methods quantitatively on our dataset in Tab.˜1 and qualitatively in Fig.˜5 (additional visualizations in Supplementary Sec.˜S2). We find that all methods benefit from using a 2D chromaticity coordinate instead of a 1D CCT, which confirms our motivation for this work. Additionally, our 2D CST-MLP surpasses traditional interpolation and nearest-neighbor methods, providing a strong solution that integrates seamlessly into the existing DNG [AdobeDNG] pipeline, which relies on linear transformations.
@@ 4.2 Lightbox Dataset Results
In-the-wild results. After calibrating the CST prediction methods with lightbox images, we evaluate on a smaller set of 150 in-the-wild photos captured with both smartphones (Pixel and Samsung). As shown in Fig.˜6, methods that rely on 2D chromaticity deliver qualitatively superior results, enabling more accurate color correction under challenging LED illumination across diverse regions containing skin tones, posters, and craft materials. We also observe that EXPINV [expinv], a per-pixel prediction method, exhibits color casts when applied to full-size images. Additional comparisons are provided in Supplementary Sec.˜S2.
Comparing CST-MLP with per-pixel MLPs. We find our 2D CST-MLP outperforms 1D MLP-based direct prediction baselines, CCCNN [cccnn] and EXPINV [expinv]. Additionally, highlighting the strength of utilizing 2D chromaticity input, our 2D chromaticity-based variants of CCCNN [cccnn] and EXPINV [expinv] achieve lower $\Delta E_{2000}$ in many cases and surpass CST-MLP under challenging illumination conditions (higher percentiles). This performance gain arises because per-pixel MLPs can model more complex mappings that depend on both input color and white point, whereas our method predicts a global transformation that depends only on the white point. However, per-pixel methods remain unsuitable for two reasons. First, CST-MLP has a footprint of 1.5 KB and 0.01 MMACs, compared to 69 KB and 13 MMACs for per-pixel methods. CST-MLP is efficient because it predicts a single global transform. In Fig.˜7, we show that the computational cost of a per-pixel MLP far exceeds our CST-MLP at various input resolutions. Second, as shown in Fig.˜6, applying per-pixel MLP methods to full-size images can introduce color casts, likely because each pixel is transformed independently, leading to spatial inconsistencies. In contrast, CST-MLP applies a single transformation across the image, ensuring spatial consistency.
@@
Figure 7: Computational cost comparison. We plot the multiply–accumulate operations (MACs, log scale) at different resolutions for 2-CST [AdobeDNG] interpolation, EXPINV [expinv], and our 2D CST-MLP. Our CST-MLP has nearly the same computational cost as traditional methods, since the CST is computed once and applied to the entire image. In contrast, approaches such as EXPINV must evaluate an MLP at every pixel.
Table 2: Ablation of CST-MLP. We evaluate CST-MLP on Pixel camera images from both the lightbox and in-the-wild scenes. We test 1D and 2D input coordinates with and without augmentation. Our results show that using xy chromaticity yields the lowest angular error among all white-point parameterizations. Finally, we extend our analysis to higher-dimensional mappings using the xy chromaticity white point.
Input coord. Mapping Noise Aug. Lightbox In-the-wild
@@
1D (CCT) Linear $(3\times 3)$ 4.11 8.48 3.68 7.05
2D (raw) Linear $(3\times 3)$ 3.58 7.99 4.02 7.81
Linear $(3\times 3)$ 3.75 8.05 4.14 8.33
2D (xy) Linear $(3\times 3)$ 3.53 7.92 3.64 7.33
Linear $(3\times 3)$ 3.60 7.95 3.32 6.71
Polynomial $(3\times 9)$ 3.67 8.78 3.67 8.78
Polynomial $(3\times 19)$ 4.13 11.37 4.13 11.37
Root Polynomial $(3\times 6)$ 3.13 7.45 3.40 7.15
Root Polynomial $(3\times 6)$ 3.27 7.73 3.07 6.77
Root Polynomial $(3\times 13)$ 3.29 7.79 3.32 7.10
Ablation of CST-MLP. In Tab.˜2, we ablate three design choices. First, we compare specifying the whitepoint in 2D xy chromaticity, which yields better performance than using 1D CCT space or 2D raw space. Second, we find that our noise augmentation strategy is crucial for our method to generalize to out-of-domain illuminant chromaticities. Finally, we compare the performance of predicting different transforms and find that linear and root-polynomial mappings yield the best results. We further ablate CST-MLP with varying layer and node counts in Supplementary Sec.˜S1.3.
Sensitivity to imperfect input white-point. Our experiments assume that the ground-truth illuminant is known to isolate the performance of various CST prediction methods. In practice, however, this assumption does not hold as white-balance estimation [ccmnet, Afifi2025] modules are imperfect. To better assess real-world performance with imperfect white-points, we evaluate our method (trained with ground truth white-points) alongside several baselines on the lightbox dataset with controlled white-point offsets (Tab.˜3). For each chart, an offset of $0\degree$, $1\degree$, $2\degree$, $3\degree$ or $10\degree$ was applied in random directions. Our results show CST-MLP achieves the lowest mean angular error when the white-point is offset by $0\degree$$2\degree$, which falls within the typical range of white-point estimators such as CCMNet [ccmnet] and Time-Aware White Balance [Afifi2025] which achieve $1.01\degree$$2.23\degree$ mean and $0.60\degree$$1.71\degree$ median angular error on datasets that include LED illuminants. We note that when the white-point offset is large ($3\degree$$10\degree$), our 2D CST-MLP performs similarly to 1D CST-MLP, but still outperforms 2-CST traditional interpolation.
Table 3: White-point error ablation. We evaluate the performance of 2-CST, 3-CST, CCCNN, EXPINV, NN, and CST-MLP with chromaticity inputs on varying degrees of white-point offsets on the Pixel lightbox dataset. The offset columns show the angular offset added to each white point during testing. White-points are shifted by sampling a random direction and applying the corresponding offset. The results show that our CST MLP is the best-performing method in the $0\degree$$2\degree$ error range, and that 1D and 2D methods are comparable under large white-point errors ($\geq 3\degree$). Methods using 2D chromaticity inputs are highlighted in blue and we again find the 2D methods to be better or comparable to 1D parametrization.
White-point offset:
Method Dim 10°
Ang. $\boldsymbol{\Delta E}$ Ang. $\boldsymbol{\Delta E}$ Ang. $\boldsymbol{\Delta E}$ Ang. $\boldsymbol{\Delta E}$ Ang. $\boldsymbol{\Delta E}$
2-CST 1D 4.56 9.18 5.08 9.80 6.28 11.30 7.87 13.16 21.05 26.87
3-CST 1D 4.53 9.16 5.05 9.78 6.25 11.27 7.84 13.13 20.97 26.80
NN [nearestneighbors] 1D 5.24 9.53 6.07 10.75 6.65 11.72 7.90 13.44 20.27 26.99
CCCNN [cccnn] 1D 4.39 8.36 4.93 9.10 5.91 10.62 7.13 12.31 17.21 23.84
EXPINV [expinv] 1D 4.31 8.09 4.88 9.05 5.98 10.67 7.32 12.42 17.29 23.80
CST-MLP 1D 4.11 8.48 4.61 9.17 5.86 10.86 7.46 12.82 19.88 25.93
CST-MLP (RP) 1D 3.77 8.19 4.42 9.00 5.80 10.72 7.52 12.71 18.50 24.91
NN [nearestneighbors] 2D 4.30 8.74 5.27 9.94 7.03 12.03 9.15 14.33 22.68 29.08
CCCNN [cccnn] 2D 4.00 8.07 4.68 8.86 5.86 10.41 7.27 12.20 17.44 23.94
EXPINV [expinv] 2D 3.84 7.77 4.62 8.70 5.95 10.40 7.40 12.23 17.31 23.67
CST-MLP 2D 3.60 7.95 4.30 8.82 5.76 10.61 7.54 12.64 20.48 25.96
CST-MLP (RP) 2D 3.27 7.74 4.03 8.59 5.59 10.39 7.42 12.42 18.31 24.36
@@
Figure 8: Colorimetric mapping results on multi-illuminant data from LSMI [lsmi]. We visualize two multi-illuminant scenes from the Galaxy camera. Row structure for each half: Top row displays raw-RGB image, white-balanced raw-RGB image, and display-referred renderings; middle row shows contrast-stretched illumination map, first raw-RGB color chart, and display-referred renderings; bottom row presents blending map, second raw-RGB color chart, and display-referred renderings. Methods from left to right: 2-CST [AdobeDNG], 1D NN [nearestneighbors], 1D CST-MLP, 2D NN [nearestneighbors], 2D CST-MLP, and Oracle. Angular errors are displayed at the bottom left of each color chart image. Key findings: (1) Our 2D chromaticity-based methods can be extended to multi-illuminant scenes with knowledge of the illumination map and per-pixel blending maps. (2) Our 2D methods continue to exceed the performance of 1D counterparts.
Table 4: LSMI results under single and mixed-illuminant settings. Comparison of 2D chromaticity-based CST-MLP and NN [nearestneighbors] (highlighted blue) with baselines on Galaxy and Sony data.
Single-illuminant Multi-illuminant
Galaxy Sony Galaxy Sony
Method Dim Ang. $\Delta E$ Ang. $\Delta E$ Ang. $\Delta E$ Ang. $\Delta E$
Oracle 1D 2.38 4.90 1.59 3.52 2.06 3.92 1.44 3.01
2-CST [AdobeDNG] 1D 5.00 8.43 8.40 9.80 3.88 6.87 8.14 9.08
NN [nearestneighbors] 1D 4.46 7.40 3.07 5.19 3.23 5.42 2.43 4.26
CST-MLP 1D 3.77 6.30 2.70 4.55 2.52 4.28 1.88 3.40
CST-MLP (RP) 1D 3.60 6.54 2.37 4.45 2.35 4.62 1.74 3.55
NN [nearestneighbors] 2D 3.98 6.56 2.48 4.39 2.83 4.76 1.93 3.61
CST-MLP 2D 3.49 5.82 2.49 4.22 2.26 3.88 1.82 3.28
CST-MLP (RP) 2D 3.37 6.12 2.31 4.50 2.14 4.20 1.69 3.63
@@ 4.3 LSMI Dataset Results
We evaluate CST-MLP on single-illuminant scenes and multi-illuminant scenes, both sourced from the LSMI [lsmi] dataset.
Single-illuminant results. We train our CST-MLP and all baseline models on single-illuminant images from the LSMI dataset, since we do not have access to lightbox captures. We filter the LSMI dataset to include scenes where all patches on all color charts are not clipped, resulting in 1504 images for Galaxy and 1414 for Sony. We then split the data into train, validation, and test sets using 50%, 20%, and 30% of the images, respectively. For this method, we are unable to compare with CCCNN [cccnn] and EXPINV [expinv] because these methods require access to the cameras’ spectral sensitivities, which are not provided with the LSMI dataset. Summarized results comparing our 2D CST-MLP with 2-CST interpolation (using DNG CST matrices), nearest-neighbor (NN), and 1D CST-MLP baselines are presented in Tab.˜4 (left). We find that our 2D CST-MLP achieves the highest performance on this dataset. Additional details and results are provided in Supplementary Sec.˜S3.
Multi-illuminant results. Assuming access to per-pixel illuminant information and blending maps that quantify each illuminant’s contribution, we can apply our MLP trained on single-illuminant images to multi-illuminant scenes. Specifically, for each illuminant, we predict and apply the CST to the white-balanced image to generate multiple corrected versions. We then merge these corrected images using the blending map as weights. A detailed description of this procedure is provided in Supplementary Sec.˜S3. In Tab.˜4 (right), we present quantitative comparisons of our method. Additionally, in  Fig.˜8, we present qualitative results on 2- and 3-illuminant scenes from the LSMI dataset. We observe that our 2D CST-MLP outperforms other methods quantitatively and qualitatively, demonstrating the benefits of using a 2D chromaticity representation in multi-illuminant scenarios. Additional details and results are provided in Supplementary Sec.˜S3.